OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 9x + y = 0 18x + 4y = –8

OpenStudy (anonymous):

okay so from 1st equation y=-9x put this value of y=-9x in second equation 18x + 4(-9x)=-8 18x-36x=-8 -18x=-8 x=8/18

OpenStudy (anonymous):

thank u